### v2

# MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in Large Language Models and Implications for AI in Education

Abstract
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of Large Language Models (LLMs) through a pedagogical approach. The dataset comprises mathematics and reading comprehension questions, each accompanied by four answer choices and their corresponding rationales. At the heart of MalAlgoQA are “malgorithms” - rationales behind incorrect answer choices that represent flawed yet logically coherent reasoning paths. These malgorithms serve as counterfactual scenarios, allowing us to assess an LLM’s ability to identify and analyze flawed reasoning patterns. We propose the Malgorithm Identification task, where LLMs are assessed based on their ability to identify corresponding malgorithm given an incorrect answer choice. To evaluate the model performance, we introduce two metrics: Algorithm Identification Accuracy (AIA) for correct answer rationale identification, and Malgorithm Identification Accuracy (MIA) for incorrect answer rationale identification. Our experiments reveal that state-of-the-art LLMs exhibit significant performance drops in MIA compared to AIA, highlighting the challenges in counterfactual reasoning. Surprisingly, we find that the chain-of-thought prompting technique not only fails to consistently enhance MIA but can sometimes lead to underperformance compared to simple prompting. These findings have important implications for developing LLMs with improved counterfactual reasoning, particularly relevant for AI-powered tutoring systems, where identifying and addressing student misconceptions is essential. MalAlgoQA dataset is available here.
MalAlgoQA: Pedagogical Evaluation of Counterfactual Reasoning in
Large Language Models and Implications for AI in Education
Naiming Liu Rice University nl35@rice.edu                        Shashank Sonkar Rice University ss164@rice.edu
MyCo Le Rice University msl12@rice.edu                        Richard G. Baraniuk Rice University richb@rice.edu
footnotetext: Equal contribution.
1 Introduction
Evaluating the counterfactual reasoning capabilities of Large Language Models (LLMs) – their ability to consider “what if” scenarios and infer causes based on hypothetical situations – remains a significant challenge despite their impressive performance in various natural language tasks (Pearl, 2019; Rafetseder et al., 2021; Sonkar and Baraniuk, 2023). We introduce a novel method to evaluate these capabilities, drawing inspiration from educational assessment techniques. Specifically, we leverage the practice of distractor choice generation, which involves creating plausible but incorrect answer options by envisioning hypothetical scenarios and logical yet flawed reasoning paths. This approach mimics how educators anticipate and model potential student misconceptions, engaging in counterfactual thinking about alternative cognitive processes that could lead to specific errors.
Question Each side of a square is 3 inches long. A student increases the length of each side by 3 inches. The area of the new square is how many times the area of the original square?
Answer Choices (A) 9 (Rationale: Squared the change in the length.) (B) 2 (Rationale: Calculated the relationship between the sides.) (C) 3 (Rationale: Used the change in the length.) (D) 4 (Rationale: Calculated the area of the original square: $3\times 3=9$ and the area of the new square with each side length increased by 3 inches: $6\times 6=36$. Then determined that the area of the new square is 4 times the area of the original square.)
Table 1: An example question from the MalAlgoQA dataset illustrating the Malgorithm Identification task. Each answer choice is associated with a rationale, representing the reasoning process that led to that answer. The task for the model is to correctly identify the rationale given a particular answer choice. For incorrect answer choices, the corresponding rationales are malgorithms, representing flawed reasoning processes and thereby evaluating the counterfactual reasoning abilities of LLMs.
Building on this foundation, we present MalAlgoQA, a dataset designed to challenge LLMs’ ability to reason about hypothetical situations and identify flawed logical pathways. MalAlgoQA consists of a diverse set of 807 mathematics and 290 reading comprehension questions spanning various grade levels (3-11), content classifications (e.g., algebra, geometry, number & operations), and Depth of Knowledge levels (DOK) (1-3) (Webb, 2002). Each question is accompanied by a set of answer choices and their associated rationales.
Central to our approach is the concept of “malgorithms” - a term we coin to describe the flawed reasoning paths that lead to incorrect answers (name is inspired by the concept of “mal-rules” proposed by Payne and Squibb (1990)). These malgorithms represent deterministic yet erroneous thought processes, mirroring the way students might logically follow a series of steps based on a misunderstanding or misapplication of concepts. We propose the Malgorithm Identification task to assess LLMs’ ability to reason about both correct and incorrect answer rationales. In this task, given a question and a specific answer choice, the model must identify the underlying rationale (algorithm or malgorithm) that led to that choice. This approach challenges LLMs to engage in both causal and counterfactual reasoning. To illustrate, consider a simple question: “What is 1 + 2 × 3 + 4?” with choices A. 11, B. 13, and C. 21. The rationales are: “Worked left to right” (corresponding to choice B), “Applied PEASMD rule instead of PEMDAS” (corresponding to choice C), and “Applied PEMDAS rule” (corresponding to the correct choice A). Given an answer choice, the model should identify the rationale explaining that choice.
In order to evaluate model performance, we introduce two key metrics: Algorithm Identification Accuracy (AIA), which measures accuracy in identifying the rationale behind a given correct answer. Malgorithm Identification Accuracy (MIA), of particular interest, evaluates the model’s counterfactual causal reasoning abilities by quantifying accuracy in identifying the rationale behind an incorrect answer. Our experiments reveal that GPT-4o (OpenAI, 2023) achieves an AIA of $95.7\%$, demonstrating its proficiency in this task. However, when it comes to MIA, GPT-4o’s performance drops significantly, with an accuracy of only $66.1\%$. This pattern, where accuracy is higher for correct answer rationales and lower for incorrect answer rationales, is consistent across other language models, such as LLaMA-3-70B (Dubey et al., 2024) and GPT-3.5 (Schulman et al., 2022). It emphasizes the challenge of counterfactual reasoning for LLMs, especially on finding the underlying rationale behind incorrect answer choices. Our experiments also reveal an unexpected result that Chain-of-Thought (CoT) prompting Wei et al. (2022b); Kojima et al. (2022) not only fails to consistently enhance MIA, it sometimes even underperforms compared to simple prompting.
The implications of our findings reveal critical challenges for AI in education. The stark performance gap between AIA and MIA across all models (e.g., GPT-4o’s $95.7\%$ AIA vs. $66.1\%$ MIA) exposes fundamental limitations in LLMs’ ability to understand student misconceptions. This discrepancy suggests that current LLM tutoring systems fail to identify and address the root causes of student errors. Such limitations necessitate a rethinking of feedback mechanisms in LLM-powered educational tools. Their struggle with malgorithm identification (e.g., GPT-3.5’s dramatic drop from $86.3\%$ AIA to $14.74\%$ MIA in Math) indicates they may offer incomplete or potentially misleading guidance when addressing student mistakes. These insights have profound implications for automated grading and assessment. The significant performance drop in MIA across all models raises concerns about the reliability of AI-driven systems in evaluating open-ended questions or complex problem-solving tasks.
Our work makes two key contributions to the field of LLMs and AI for education, with critical implications for developing more effective AI-powered tutoring systems. First, we introduce MalAlgoQA in section 2, a novel dataset inspired by pedagogical distractor generation techniques, specifically designed to evaluate the counterfactual reasoning capabilities of LLMs. This approach mimics how educators create plausible but incorrect answer options, challenging LLMs to engage in the type of reasoning essential for understanding and addressing student misconceptions. Second, we propose the Malgorithm Identification task in section 3, along with two associated evaluation metrics: AIA and MIA. Our findings in sections 4 and 5 reveal a significant performance gap between these metrics across various LLMs, highlighting the challenge these models face in counterfactual reasoning. We discuss the implications of our findings for AI-powered educational tool in section 6. By developing methods to evaluate LLMs’ ability to identify and understand student misconceptions, our work paves the way for more effective AI tutoring systems that can accurately diagnose and address individual students’ conceptual gaps Liu et al. (2023); Sonkar et al. (2023, 2024c).
2 MalAlgoQA Dataset
We introduce MalAlgoQA, a real-world dataset consisting of multiple-choice questions in mathematics and reading comprehension for students in grade 3 through 11. The dataset is designed to evaluate the counterfactual reasoning abilities of LLMs through the task of analyzing the rationales and reasons behind incorrect answer choices, a process we refer to as “Malgorithm identification.” The MalAlgoQA dataset provides valuable resources to assess the counterfactual reasoning capabilities of LLMs through a pedagogical approach.
2.1 Dataset Overview
MalAlgoQA comprises two multiple-choice question sets: a mathematics and a reading comprehension question set. The mathematics question set contains 807 questions, while the reading comprehension question set has 290 questions. Each question is associated with four answer choices, as well as their corresponding rationales. Reading question set also provides a passage for each question. Table 1 provides an example of math question, with more examples available in A. Additionally, the dataset provides rich metadata including the grade level, content classification, and Depth of Knowledge (DOK) level for each question.
2.1.1 Grade Levels
The questions in MalAlgoQA span across multiple grade levels, ranging from Grade 3 to Grade 11. As illustrated in Table 2, the distribution of questions for both mathematics and reading comprehension subsets is balanced across different grade levels.
Grade Level Math Reading
3 132 45
4 120 47
5 94 33
6 140 41
7 159 35
8 128 42
10 - 47
11 34 -
Table 2: Distribution of questions across grade levels.
2.1.2 Content Classifications
The mathematics questions set are categorized into five content classifications: Number & Operation, Algebra, Geometry & Measurement, Data Analysis, and Data Analysis & Probability. The reading comprehension questions are divided into two content classifications: Informational Text and Literature. Table 3 presents an overview of the distribution across content classifications for both question sets.
Math Content Classifications
Classification Questions
Algebra 315
Number & Operation 312
Geometry & Measurement 122
Probability 44
Data Analysis 14
Reading Content Classifications
Classification Questions
Informational Text 190
Literature 100
Table 3: Distribution of questions across content classifications.
2.1.3 Depth of Knowledge (DOK) Levels
Each question in MalAlgoQA is annotated with a DOK level, which indicates the cognitive complexity of the question. The DOK levels are derived from Norman L. Webb’s taxonomy Webb (2002) and range from I to III, with higher levels representing increased cognitive complexity. Table 4 shows the distribution of questions across DOK levels for both question sets.
DOK Level Math Reading
I 349 37
II 434 203
III 24 50
Table 4: Distribution of questions across DOK levels.
2.2 Question and Rationale Characteristics
The questions in MalAlgoQA vary in length and complexity. The average length of mathematics questions is 122 characters, while the average length of reading comprehension questions is 104 characters. Reading comprehension questions are accompanied by passages, which have an average length of 5783 characters. Each answer choice is associated with an rationale, which provide explanations for the correct and incorrect answer choices. The average length of answer rationales for mathematics questions is 227 characters, while the average length of answer rationales for reading comprehension questions is 1699 characters.
3 Methodology
CoT Prompt for Malgorithm Identification:
System: Provide step-by-step reasoning to determine the answer rationale that corresponds to the given choice for the question. Then, provide your answer in the specified JSON format.
Question: {question}
Choice: {choice}
Answer Rationales: {formatted_rationales}
Given the question and answer rationales, determine which answer rationale corresponds to the given choice for the question.
Provide step-by-step reasoning. Show the steps.
And then provide your answer in the following JSON format:
"Correct Choice": "[A/B/C/D]"
Table 5: Prompt for the Malgorithm Identification experiment, where the LLM identifies the relationship between answer choices and rationales in a multiple-choice question-answering setting. The task involves identifying the “malgorithm”, a rationale that represents flawed reasoning steps which lead to an incorrect answer choice. The prompt provides the question, an answer choice (either correct or incorrect), and all rationales, asking the LLM to identify the rationale corresponding to the given choice. The highlighted sentence instructs the model to provide step-by-step reasoning, an instruction added for the Chain-of-Thought prompting strategy.
In this section, we present the task formulation for Malgorithm Identification.
3.1 Definitions and Task Formulation
Let $Q=\{q_{1},q_{2},...,q_{n}\}$ be a set of multiple-choice questions, where each question $q_{i}$ has a set of answer choices $C_{i}=\{c_{i1},c_{i2},c_{i3},c_{i4}\}$ and a set of rationales $R_{i}=\{r_{i1},r_{i2},r_{i3},r_{i4}\}$ corresponding to each answer choice. The rationale can be thought of as a chain of reasoning steps, where each step represents a part of the thought process leading to the selection of an answer choice.
Malgorithm
We introduce the term “malgorithm” to describe a rationale underlying an incorrect answer choice. A malgorithm is composed of a series of reasoning steps, some of which are correct, while others contain errors. We refer to these erroneous steps as “mal-rules”. The term mal-rules is derived from the cognitive error literature (Brown and VanLehn, 1980; VanLehn, 1990; Payne and Squibb, 1990), which refers to the flawed or incorrect rules that students may apply when solving problems, resulting in systematic errors. In our context, a malgorithm is a reasoning chain that consists of both correct rules and mal-rules, with the presence of mal-rules ultimately leading to the selection of an incorrect answer choice.
Malgorithm Identification Task
Given a question $q_{i}$, an answer choice $c_{ij}$ (either correct or incorrect), and the set of rationales $R_{i}$, the LLM must identify the rationale $r_{ik}$ that corresponds to the given answer choice $c_{ij}$. If the given answer choice is incorrect, the corresponding rationale is a malgorithm containing a combination of correct reasoning steps and one or more mal-rules.
3.2 Evaluation Metrics
To gain a comprehensive understanding of the LLMs performance in the Malgorithm Identification task, we use two key metrics:
1. Algorithm Identification Accuracy (AIA): This metric measures the model’s ability to identify the rationale or algorithm that corresponds to the correct answer choice. Essentially, AIA evaluates how well the model recognizes and follows a chain of reasoning steps that lead to the correct answer.
2. Malgorithm Identification Accuracy (MIA): This metric measures the model’s ability to identify the flawed reasoning or malgorithm that corresponds to an incorrect answer choice. Specifically, MIA evaluates the model’s capacity to identify faulty reasoning chains that contain a combination of correct steps and one or more mal-rules, which lead to the wrong answer.
4 Experiments and Findings
LLM Subject Num Shots Prompting MCQ AIA MIA
GPT 4o Reading 0 Simple 95.86 98.23 82.80
0 CoT 95.52 97.88 68.63
LLaMA3-70B Reading 0 Simple 89.45 95.72 60.00
0 CoT 89.27 94.60 58.40
GPT-3.5 Reading 0 Simple 84.83 95.05 13.98
0 CoT 81.66 85.82 8.61
LLaMA3-8B Reading 0 Simple 73.96 89.62 19.89
0 CoT 78.05 89.52 15.47
GPT 4o Math 0 Simple 90.11 92.06 62.82
5 Simple 90.61 64.99
0 CoT 97.61 95.65 66.10
5 CoT 95.49 65.10
LLaMA3-70B Math 0 Simple 73.83 94.31 50.22
5 Simple 93.65 48.31
0 CoT 90.72 95.47 55.32
5 CoT 95.89 50.6
GPT 3.5 Math 0 Simple 52.85 77.08 35.8
5 Simple 76.85 34.61
0 CoT 82.53 86.28 14.74
5 CoT 81.87 17.06
LLaMA3-8B Math 0 Simple 45.79 75.29 34.81
5 Simple 80.32 38.41
0 CoT 70.68 83.07 27.05
5 CoT 83.05 31.39
Table 6: Performance results of LLMs on the Malgorithm Identification task for Reading and Math subjects. Results are organized by prompting type (Simple/CoT) and number of shots (0/5). Metrics include Multiple-Choice Question (MCQ) accuracy, Algorithm Identification Accuracy (AIA), and Malgorithm Identification Accuracy (MIA). AIA measures correct rationale identification, while MIA measures flawed reasoning identification. MIA scores are generally significantly lower than AIA, highlighting the challenge of counterfactual reasoning.
Subject Grade GPT 4o LLaMA3-70B
Math 3 67.00 44.26
4 65.93 54.13
5 66.52 46.97
6 67.80 58.84
7 64.50 57.49
8 62.96 42.44
11 53.68 41.03
average 65.10 50.60
Subject Grade GPT 4o LLaMA3-70B
Reading 3 73.81 65.18
4 68.89 60.71
5 69.47 45.68
6 58.62 50.00
7 63.46 80.22
8 80.67 62.00
10 64.96 46.83
average 68.63 58.40
Table 7: Grade-level performance distribution (MIA) for GPT-4o and LLaMA3-70B models using Chain-of-Thought prompting (few-shot for Math, zero-shot for Reading). Grades with above-average performance are underlined.
4.1 Experimental Setup
We conducted experiments on the Malgorithm Identification task with four state-of-the-art LLMs: GPT-4o, GPT-3.5, LLaMA3-70B and LLaMA3-8B.The specific models we use are GPT-4o: gpt-4o-2024-05-13, GPT-3.5: gpt-3.5-turbo-0125, LLaMA3-70B: Meta-LLaMA3-70B-Instruct, LLaMA3-8B: Meta-LLaMA3-8B-Instruct To evaluate the LLMs performance, we use AIA and MIA metrics (described in section 3.2).
For each model, we evaluated with four experimental settings: (1) Simple prompting, where the LLMs are prompted to provide only the final answer; (2) Chain-of-Thought (CoT) prompting Wei et al. (2022b); Kojima et al. (2022), where the LLMs are prompted to provide step-by-step reasoning and show the intermediate steps before giving the final answer, as shown in Table 5; (3) Zero-shot learning, where the LLMs are provided without any additional examples; (4) Few-shot learning, where the LLMs are provided with a small number of examples demonstrating the desired reasoning process, in addition to the standard prompts. Due to context length constraints, we conducted few-shot experiments only on the Math dataset.
4.2 Main Findings
Wide Performance Gap Between MIA and AIA
The most striking observation from Table 6 is the substantial performance gap between Malgorithm Identification Accuracy (MIA) and Algorithm Identification Accuracy (AIA) across all models. For GPT-4o on the Math dataset with zero-shot CoT prompting, we see an AIA of $95.65\%$ compared to an MIA of only $66.10\%$ a drop of nearly $30$ percentage points. This pattern is consistent across models and prompting strategies, with the gap widening for smaller models. For instance, GPT-3.5 shows an even more dramatic difference: $86.28\%$ AIA vs $14.74\%$ MIA.
This disparity suggests that while LLMs have become proficient at identifying correct reasoning paths, they struggle significantly when asked to recognize and articulate flawed reasoning. This challenge goes beyond simple counterfactual reasoning and points to a fundamental limitation in how these models process and evaluate logical structures. The ability to identify correct reasoning does not necessarily translate to the ability to pinpoint errors in incorrect reasoning, indicating a potential blind spot in the training or architecture of current LLMs.
Impact of Model Size on MIA Performance
The data reveals a clear correlation between model size and MIA performance. For the Math dataset with zero-shot CoT prompting, we see a progression from LLaMA3-8B ($27.05\%$ MIA) to LLaMA3-70B ($55.32\%$ MIA) to GPT-4o ($66.10\%$ MIA). This trend suggests that the ability to identify malgorithms is an emergent property that significantly improves with increased model capacity Wei et al. (2022a). However, it’s important to note that this improvement with model size is much more pronounced for MIA than for AIA. While GPT-4o outperforms LLaMA3-8B by $39.05$ percentage points in MIA, the difference in AIA is only $12.58$ percentage points ($95.65\%$ vs $83.07\%$). This disproportionate scaling indicates that malgorithm identification may require more complex reasoning capabilities that are only unlocked at larger model scales.
Figure 1: An illustration of MIA performance across different Grades and DOK levels for Math and Reading question sets. The top number represents MIA performance and the bottom number represents question counts. DOK I and II are selected for Math and DOK II and III for Reading. (best viewed in colors).
Ineffectiveness of Few-Shot Learning
The data clearly demonstrates that few-shot learning does not significantly improve MIA performance across models. For GPT-4o on the Math dataset, the MIA score with CoT prompting actually decreases slightly from $66.10\%$ (zero-shot) to $65.10\%$ (5-shot). LLaMA3-70B shows a similar trend, with MIA dropping from $55.32\%$ to $50.6\%$. Even in cases where there is a marginal improvement, the gains are negligible. For instance, LLaMA3-8B’s MIA increases from $27.05\%$ to $31.39\%$ with 5-shot CoT prompting - a difference of only $4.34$ percentage points. GPT-3.5 shows a similarly small improvement from $14.74\%$ to $17.06\%$.
These results suggest that providing examples does not substantially enhance an LLM’s ability to identify malgorithms. This finding is particularly noteworthy given that few-shot learning often improves performance on other NLP tasks. The ineffectiveness of this approach for malgorithm identification underscores the unique challenges posed by this task and suggests that more sophisticated techniques may be needed to improve LLM performance in this area.
4.3 Understanding CoT Prompting for MIA
Our analysis of Chain-of-Thought (CoT) prompting reveals an unexpected ineffectiveness in improving Malgorithm Identification Accuracy (MIA), and in some cases, even leads to decreased performance compared to simple prompting.
Inconsistent impact across subjects: For GPT-4o, CoT prompting shows negligible difference in Math MIA ($65.1\%$ vs $65.0\%$ for simple prompting), but significantly underperforms in Reading MIA ($68.6\%$ vs $82.8\%$ for simple prompting).
Consistent underperformance of LLaMA-70B: LLaMA-70B demonstrates lower MIA prompting in both Reading ($58.4\%$ vs $60.0\%$) and Math ($48.3\%$ vs $50.6\%$) with CoT compared to simple prompting.
Minimal impact on AIA: CoT prompting generally performs similarly to simple prompting for Algorithm Identification Accuracy (AIA) across models and subjects, indicating its ineffectiveness is specific to malgorithm identification.
Figure 2: An illustration of MIA performance across five math content classes (Number & Operation, Algebra, Geometry & Measurement, Data Analysis, Probability) and two Reading content classes (Informational, Literature).
These results challenge the assumption that explicit step-by-step reasoning always improves performance on complex tasks. The data suggests that CoT prompting, while effective for many NLP tasks, may actually hinder the model’s ability to identify flawed reasoning paths in the context of malgorithm identification.
A plausible explanation for this counterintuitive finding lies in the nature of the training data typically used for LLMs. These models are predominantly exposed to examples of correct reasoning during pre-training and fine-tuning. Consequently, when CoT prompting guides the model through a structured thought process, it likely leverages these learned patterns of correct reasoning. This may inadvertently anchor the model to correct logical pathways, making it more challenging to deviate from these patterns and identify flawed reasoning. In essence, CoT prompting might be reinforcing the model’s bias towards correct reasoning, which is counterproductive when the task specifically requires identifying incorrect logical steps.
5 Analysis: Educational Dimensions
Performance across Grade Levels
Table 7 analyzes the MIA performance of GPT-4o and LLaMA3-70B models across different grade levels in math and reading datasets. Both models consistently show a decline in performance with increasing grade levels, with LLaMA3-70B exhibiting some unexpected variability, particularly excelling in Grade 8 of the reading dataset. These results highlight the increasing challenges in tackling malgorithm identification task as the questions becomes more complex.
Performance across DOK
Figure 1 presents a comprehensive overview of performance across DOK levels for various Grade using MIA score from GPT 4o with CoT prompting strategy. The pattern indicates that as the DOK level and Grade increases, MIA performance tends to decline, especially highlighted by the comparison between the top-left cell (dark green) and bottom-right cell (dark red). Similarly, in the Reading dataset, performance at DOK Level III drops to 33.33 for Grade 7. This trend further emphasizes the inherent difficulty in evaluating the malgorithms of questions that require deeper understanding and analytical thinking.
Performance across Content Classifications
Figure 2 illustrates the performance of LLMs in detecting malgorithms across five math contents and two reading content classifications. For the math dataset, Geometry consistently yields the highest performance across all models, suggesting that even for malgorithm identification, structural content like Geometry is easier for LLMs to handle as compared to numerical content. On the other hand, Probability appears to be more challenging content, reflecting the inherent difficulty of LLMs in probabilistic reasoning. For the reading dataset, GPT-4o and LLaMA3-70B excel at Informational Text, while smaller models like GPT-3.5 and LLaMA3-8B perform better with Literature. This observation suggests that larger models are better at processing structured, factual content, whereas smaller models are more efficient with narrative, context-rich material.
6 Implications for AI in Education
The findings from our study on malgorithm identification have significant implications for the development and application of AI in educational contexts:
Limitations in Understanding Student Misconceptions: The substantial performance gap between Algorithm Identification Accuracy (AIA) and Malgorithm Identification Accuracy (MIA) suggests that current LLMs are far more adept at recognizing correct reasoning than identifying flawed logic. For GPT-4o, we observed a drop from $95.65\%$ AIA to $66.10\%$ MIA on the Math dataset. This disparity indicates that AI tutoring systems based on these models may excel at confirming correct answers but struggle to diagnose and address student misconceptions effectively. This limitation could lead to missed opportunities for targeted intervention and personalized learning.
Rethinking Training Approaches for Error Recognition The ineffectiveness of CoT prompting in improving MIA performance reveals a fundamental challenge in training LLMs to recognize and analyze errors. Rather than simply diversifying training data, this finding suggests a need for novel training paradigms that specifically target the skill of error identification Sonkar et al. (2024a). Future AIED systems may require specialized fine-tuning processes or architectural modifications that enable models to simultaneously reason about correct and incorrect logical pathways. This could involve developing new pre-training tasks focused on error detection, or creating model architectures that maintain separate representations for correct and flawed reasoning patterns. Such innovations could lead to AI tutors that are equally adept at reinforcing correct understanding and diagnosing misconceptions, significantly enhancing their effectiveness in personalized education.
Rethinking Feedback Mechanisms: The challenges in malgorithm identification suggest that current AI models might provide incomplete or misleading feedback to students. AIED systems need to be designed with an awareness of these limitations, potentially incorporating human oversight or alternative assessment methods to ensure that student misconceptions are accurately identified and addressed.
Implications for Automated Grading and Assessment: The performance discrepancy between AIA and MIA also raises concerns about the reliability of AI-driven automated grading systems, particularly for open-ended questions or problems requiring complex reasoning. Educational institutions considering the implementation of such systems should be aware of these limitations and consider alternate Sonkar et al. (2024b, d) or hybrid approaches with human review.
Adaptive Difficulty Handling: The observed decline in MIA performance across increasing grade levels and DOK levels highlights the need for AIED systems that can adapt to varying levels of question complexity. Future AI tutors should be designed with mechanisms to maintain consistent performance across different difficulty levels, ensuring effective support for students as they progress through more challenging material.
The difficulty LLMs face in identifying flawed reasoning points to a broader challenge in counterfactual reasoning. For AIED applications, this suggests a need to develop models with enhanced capabilities in exploring alternative scenarios and reasoning paths. Such improvements could lead to more sophisticated tutoring systems that can not only identify errors but also explain why certain approaches are incorrect and guide students towards correct understanding. By addressing these implications, AIED researchers can work towards not only improving the accuracy of student assessment but also enhancing the ability of AI systems to provide targeted and personalized support.
7 Related Work
Counterfactual reasoning, a critical cognitive process for understanding causality, has gained significant attention in machine learning (ML) and NLP research. Pearl’s seminal work on causal inference Pearl (2009) has been widely applied to address challenges in fairness Kusner et al. (2017), interpretability Wachter et al. (2018), and robustness Bareinboim and Pearl (2016) in ML. In NLP, counterfactual reasoning is important for tasks requiring comprehension of hypothetical scenarios. However, most existing NLP models primarily focus on predictive tasks rather than hypothetical reasoning Bommasani et al. (2021); Bender et al. (2021). Recent work has begun to address this gap developing specialized datasets and models for counterfactual reasoning Qin et al. (2019). Several datasets have been created to evaluate different aspects of reasoning, including counterfactual reasoning such as COPA dataset Roemmele et al. (2011), Counterfactual Story Rewriting task Qin et al. (2020), and CQA dataset Talmor et al. (2019). These efforts are complemented by development of new modeling approaches, such as causal LMs Feder et al. (2021), counterfactual data augmentation Kaushik et al. (2020), and generating counterfactual explanations Wu et al. (2021); Ross et al. (2021). Despite these advancements, challenges remain such as handling multi-step causal chains and generalizing across diverse domains Xu et al. (2023); Keith et al. (2020); Feder et al. (2022).
8 Conclusion and Future Work
This paper introduces MalAlgoQA, a novel dataset designed to evaluate the counterfactual reasoning capabilities of LLMs through a pedagogical lens. By proposing the Malgorithm Identification task and its associated metrics, we provide a new framework for assessing LLMs’ ability to recognize and analyze flawed reasoning patterns. Our findings reveal significant challenges in counterfactual reasoning for state-of-the-art LLMs. These results highlight the need for further research into enhancing LLMs’ counterfactual reasoning abilities, particularly for educational applications where understanding student misconceptions is crucial. Future work will focus on expanding the dataset, exploring novel training methodologies, and investigating architectural modifications to develop more cognitively-aligned language models capable of robust counterfactual reasoning.
9 Limitations
While our research sheds light on the counterfactual reasoning capabilities of LLMs, we acknowledge the inherent complexity and intricacy of the cognitive process may not be fully captured by our current evaluation framework. The MalAlgoQA dataset and the Malgorithm Identification task, although designed with careful consideration, are initial steps toward understanding the intricate aspect of cognition. The chain-of-thought (CoT) prompting technique is not a panacea and its effectiveness vary depending on the specific context and task. As research in this field progresses, we anticipate further refinements and enhancements to our methodology, contributing to a more comprehensive understanding of counterfactual reasoning in LLMs.
10 Ethics and Risk
The use of LLMs in classrooms, for instance, could profoundly influence the learning experiences of students. While LLMs can provide personalized learning support and instant feedback, their inability to accurately identify and address student misconceptions, as highlighted by our study, could potentially reinforce incorrect understanding or reasoning. However, the transparency of LLMs’ decision-making process is a crucial ethical consideration. If LLMs are to be trusted educational aids, it is critical for students, teachers, and parents to understand the reasoning behind the LLMs’ responses. Our work with the MalAlgoQA dataset is a significant step towards understanding and addressing these ethical considerations. By examining the counterfactual causal reasoning abilities of LLMs, we can assess their readiness for use in educational settings and contribute to the ongoing discourse surrounding the ethical deployment of AI in education. However, we acknowledge that this is a complex, multifaceted issue that requires continuous exploration, vigilance, and open dialogue among researchers, educators, policymakers, and society at large.
Acknowledgments
This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.
Appendix A Appendix
Examples of MalAlgoQA dataset
We provide five math examples, one per content classifications, of question, choices and their corresponding rationales in Table 8 and 9.
Math Question Set Example 1:
Content Classification: Number & Operation
Question: What number is subtracted from 1,000 to result in a difference of 421?
Answer and Rationales:
A: 421
Rationale A: Selects the result of subtraction.
B: 579
Rationale B: 1000 - 421 = 579
C: 621
Rationale C: Rounded the hundreds place and added the tens and ones place. 1000 - 400 = 600; 600 + 21 = 621
D: 689
Rationale D: Rounded in hundreds, tens and ones place. 1000 - 400 = 600; 100 - 20 = 80; 10 - 1 = 9; 600 + 80 + 9 = 689
Math Question Set Example 2:
Content Classification: Algebra
Question: Anne bought a calculator that cost 30. She received 10% off her purchase and then was charged 6% tax. What was the total amount that Anne paid?
Answer and Rationales:
A: 21.20
Rationale A: Subtracted 10 from 30, then added 6%.
B: 25.38
Rationale B: Subtracted 10% of the price from 30 then subtracted 6%.
C: 28.62
Rationale C: Subtracted 10% of the price from 30 and then added 6%.
D: 28.80
Rationale D: Calculated tax based on original price, then subtracted the 3.
Math Question Set Example 3:
Content Classification: Geometry and Measurement
Question: Which side lengths form a right triangle?
Answer and Rationales:
A: 2 cm, 4 cm, 8 cm
Rationale A: Multiplied the 2 smaller sides to get the longest side.
B: 4 cm, 5 cm, 6 cm
Rationale B: Selected side with constant difference between the sides.
C: 5 cm, 12 cm, 13 cm
Rationale C: 5^2 + 12^2 = 13^2
D: 9 cm, 16 cm, 25 cm
Rationale D: Added the 2 smaller sides to get the longest side.
Table 8: Examples of Mathematics question set.
Math Question Set Example 4:
Content Classification: Data Analysis
Question: Mrs. Castillo recorded the number of students in each grade 5 classroom.
24 28 25 24 29
What is the mean of the data?
Answer and Rationales:
A: 24
Rationale A: Chose the mode.
B: 25
Rationale B: Chose the median. {24, 24, 25, 28, 29}
C: 26
Rationale C: (24 + 28 + 25 + 24 + 29) / 5 = 130/5 = 26
D: 29
Rationale D: Chose the greatest number.
Math Question Set Example 5:
Content Classification: Probability
Question: Ms. Collier had a deck of cards. There were stars on 1/4 of the cards in the deck. After randomly picking a card and returning it to the deck 100 times, the expected result and relative frequency of picking a card with a star were equal. In the first 50 cards she picked, she got a star 10 times. How many times did Ms. Collier get a star in the second 50 cards she picked?
Answer and Rationales:
A: 10
Rationale A: Chose the number of times a star is picked in the 1st 50 cards drawn.
B: 13
Rationale B: Calculated 50 * 1/4 = 12.5 and then rounded to 13.
C: 15
Rationale C: Correct. 100 * 1/4 = 25 cards w/ stars. 25 - 10 = 15
D: 25
Rationale D: Calculated 100 * 1/4 = 25 cards w/ stars.
Table 9: Examples of Mathematics question set.